40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 17 23:36:12 2018, johannes, Configuration, General, AS port laser injection PSLbeat.svg.pngASpath.svg.png
    Reply  Wed Apr 18 22:46:23 2018, johannes, Configuration, General, AS port laser injection ASpath.svg.pngpower_out_fluctuation_DC.pngpower_out_fluctuation_AC_zoom.png
       Reply  Thu Apr 19 01:04:00 2018, gautam, Configuration, General, AS port laser injection 
          Reply  Thu Apr 19 20:41:09 2018, Koji, Configuration, General, Aux Laser LD dying? (AS port laser injection) Aux_laser_adj_Pout.pdfAux_laser_RIN.pdf
             Reply  Tue Apr 24 08:36:47 2018, johannes, Configuration, General, Aux Laser LD dying? (AS port laser injection) 
       Reply  Thu May 3 10:40:43 2018, johannes, Configuration, General, AS port laser injection mm_solution_AStable.pngAStable_beampath.pdf
          Reply  Thu May 3 12:10:12 2018, gautam, Configuration, General, AS port laser injection 
          Reply  Wed May 23 00:45:18 2018, johannes, Configuration, General, AS port laser injection faraday_location.pdf
Message ID: 13764     Entry time: Wed Apr 18 22:46:23 2018     In reply to: 13555     Reply to this: 13766   13810
Author: johannes 
Type: Configuration 
Category: General 
Subject: AS port laser injection 

Using Gautam's Finesse file and the cad files for the 40m optical setup I propagated the arm mode out of the AS port. For the location of the 3.04 mm waist I used the average distance to the ITMs, which is 11.321 m from the beam spot on the 2 inch mirror on the AS table close to the viewport. The 2inch lens focuses the IFO mode to a 82.6 μm waist at a distance of 81 cm, which is what we have to match the aux laser fiber output to.

I profiled the fiber output and obtained a waist of 289.4 μm at a distance of 93.3 cm from the front edge of the base of the fiber mount. Next step is to figure out the lens placement and how to merge the beam paths. We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


While doing a beam scan with the razor blade method I noticed that the aux laser has significant intensity noise. This is seen on the New Focus 1611 that is used for the beat signal between PSL and aux laser, as well as on the fiber output PD. There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. Koji suggested it could be light scattering back into the laser because I couldn't find a usable Faraday Isolator back when I installed the aux laser in the PSL enclosure. I'll have to investigate this a little further, look at the spectrum, etc. This intensity noise will appear as amplitude noise of the beat note, which worries me a little.

power_out_fluctuation_DC.png      power_out_fluctuation_AC_zoom.png

Attachment 1: ASpath.svg.png  3.687 MB  Uploaded Thu Apr 19 00:38:37 2018  | Hide | Hide all
ASpath.svg.png
ELOG V3.1.3-